Problem 5-14

The following information relates to Pukalani Industries for fiscal 2017, the company’s first year of operation:
Units produced 494,900
Units sold 448,800
Units in ending inventory 46,100
Fixed manufacturing overhead $1,410,465
Calculate the amount of fixed manufacturing overhead that would be expensed in 2017 using full costing. (Round fixed manufacturing overhead per unit to 2 decimal places, e.g. 15.25 and final answer to 0 decimal places, e.g. 125.)
Fixed manufacturing overhead expensed $

LINK TO TEXT

LINK TO TEXT

Calculate the amount of fixed manufacturing overhead that would be expensed in 2017 using variable costing.
Fixed manufacturing overhead expensed $

LINK TO TEXT

LINK TO TEXT

Calculate the amount of fixed manufacturing overhead that would be included in ending inventory under full costing and reconcile it to the difference between parts a and b.
Fixed manufacturing overhead in ending inventory $

FMOH expensed under variable costing $

FMOH expensed under full costing

Difference $
Answer #1

1) Fixed manufacturing overhead expense under full costing = 1410465/494900*448800 =1279080

2) Fixed manufacturing overhead expense under variable costing = 1410465

3) Fixed manufacturing overhead in ending inventory = 1410465/494900*46100 = 131385

FMOH expensed under variable costing = 1410465

FMOH expensed under full costing = 1279080

Difference = 131385